OpenStudy (anonymous):

y=sin^-1(lnx) differentiate with respect to x. cant we use the formula sin^-1(x)=1/root(1^2-x^2) straight away?

OpenStudy (anonymous):

yes plus chain rule go give \[\frac{1}{\sqrt{1-\ln^2(x)}}\times \frac{1}{x}\]

OpenStudy (anonymous):

but the answer is 1/xcosy

myininaya (myininaya):

\[y=\sin^{-1}(\ln(x))=>\sin(y)=\ln(x)\] \[=>\cos(y) \cdot y'=\frac{1}{x}=>y'=\frac{1}{x \cos(y)}\]

OpenStudy (zarkon):

\[\cos(y)=\sqrt{1-\sin^2(y)}\] \[\sqrt{1-\sin^2(y)}=\sqrt{1-\ln(x)^2}\] since \[\sin(y)=\ln(x)\]

OpenStudy (anonymous):

yes why do we move the sin inverse over? can't we use the sin^-1 formula straight?

OpenStudy (anonymous):

i mean why we multiply by sinx

myininaya (myininaya):

why would i remember extra formulas?

OpenStudy (anonymous):

but to do questions involving sin^-1, there's this square root formula right

OpenStudy (zarkon):

this \[\frac{d}{dx}\sin^{-1}(x)=\frac{1}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

yes!

myininaya (myininaya):

we can derive that based on what i did

OpenStudy (anonymous):

cant we apply to that question too

OpenStudy (zarkon):

that is what satellite73 did

OpenStudy (anonymous):

what did he do i don't quite understand..

OpenStudy (zarkon):

\[\frac{d}{dx}\sin^{-1}(g(x))=\frac{1}{\sqrt{1-g(x)^2}}g'(x)\]

OpenStudy (anonymous):

oh but that isn't the answer in my book...

myininaya (myininaya):

\[y=\sin^{-1}(x)=>\sin(y)=x=>\cos(y) y'=1=>y'=\frac{1}{\cos(y)}\] now if sin(y)=x/1=opp/hyp then we can find the adjacent side to y using Pythagorean thm \[adj^2=hyp^2-opp^2=1^2-x^2=>adj=\sqrt{1-x^2}\] \[y'=\frac{1}{\sqrt{1-x^2}}\]

myininaya (myininaya):

i just put y in terms of x since we began with x

OpenStudy (anonymous):

so is it that both answers are the same?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

but y prime has no x 1/x

OpenStudy (zarkon):

that is the derivative of the natural log

myininaya (myininaya):

they are same \[\cos(y)=\sqrt{1-x^2}\]

myininaya (myininaya):

well looking at when we have y=sin inverse of x

myininaya (myininaya):

what was complicated about my way?

myininaya (myininaya):

i think it is actually easier than what satellite did

OpenStudy (anonymous):

can you re explain all over again?

myininaya (myininaya):

\[y=\sin^{-1}(\ln(x))=>\sin(y)=\ln(x)\] \[\cos(y)\cdot y'=\frac{1}{x} => y'=\frac{1}{x \cos(y) }\]

myininaya (myininaya):

we can write cos(y) in terms of x like satellite has his

OpenStudy (anonymous):

so there is two answers to this question? the other one we use sin inverse x forumla?

myininaya (myininaya):

imagine a right triangle \[\sin(y)=\frac{\ln(x)}{1}=\frac{oppsite}{hypnoteuse} \] we can find the adjacent side using the Pythagorean thm \[adj^2=hyp^2-opp^2=1^2-(\ln(x))^2=>adj=\sqrt{1-(\ln(x))^2}\] so \[\cos(y)=\frac{adjacent}{hypontenues}=\frac{\sqrt{1-(\ln(x))^2}}{1}=\sqrt{1-(\ln(x))^2}\]

myininaya (myininaya):

satellite and me have same answer see

myininaya (myininaya):

\[y'=\frac{1}{x \sqrt{1-(\ln(x))^2} }\] and we don't have to remember extra formulas

OpenStudy (anonymous):

y′=1x1−(ln(x))2−−−−−−−−−−√ for this, the root 1-lnx square, why is it one?

myininaya (myininaya):

i can't read that

OpenStudy (anonymous):

y′=1x1−(ln(x))2−−−−−−−−−−√ for this, the root 1-lnx square, why is it one?

OpenStudy (anonymous):

it's satellite's answer but im just asking why is the denominator with square root, where did the one come frm?

myininaya (myininaya):

i showed you above

myininaya (myininaya):

maybe you can get zarkon to explain it better than me but that is as good as my explanation will get im sorry